OpenStudy (anonymous):

If ƒ(x ) = x^2 + 1 and g(x ) = 3x + 1, find [ƒ(4)] 2. 81 257 289

OpenStudy (anonymous):

@joyraheb

OpenStudy (anonymous):

@ilovewolf

OpenStudy (anonymous):

@Rockiee

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

???

OpenStudy (bohotness):

(gof)(x)=g(f(x)) =g(4x+2) =3(4x+2) -1 substitute 4x+2 for x in g(x) =12x +6 -1 =12x+5 Then (gof)(x)=12x+5